OpenStudy (anonymous):

What is the unit fraction of 9/10?

OpenStudy (akashdeepdeb):

divide the numerator and denominatorby 9 so that it is unity! So unit fraction should be = 1/(10/9) = 1/1.11111111111111 Getting me? :)